# Treatment Non-Adherence Bias in Clinical Machine Learning: A Real-World Study on Hypertension Medication

### v1

@@ Abstract
Machine learning systems trained on electronic health records (EHRs) increasingly guide treatment decisions, but their reliability depends on the critical assumption that patients follow the prescribed treatments recorded in EHRs. Using EHR data from 3,623 hypertension patients, we investigate how treatment non-adherence introduces implicit bias that can fundamentally distort both causal inference and predictive modeling. By extracting patient adherence information from clinical notes using a large language model, we identify 786 patients (21.7%) with medication non-adherence. We further uncover key demographic and clinical factors associated with non-adherence, as well as patient-reported reasons including side effects and difficulties obtaining refills. Our findings demonstrate that this implicit bias can not only reverse estimated treatment effects, but also degrade model performance by up to 5% while disproportionately affecting vulnerable populations by exacerbating disparities in decision outcomes and model error rates. This highlights the importance of accounting for treatment non-adherence in developing responsible and equitable clinical machine learning systems.
@@ 1 Introduction
Treatment non-adherence is a pervasive and persistent challenge in healthcare. Researchers estimate that poor medication adherence leads to 125,000 preventable deaths annually in the U.S. and contributes to \$100-\$300 billion in avoidable healthcare costs (Benjamin,, 2012). This issue is particularly prevalent among patients with chronic conditions such as hypertension, with 40-50% failing to take their medications as prescribed (Kleinsinger,, 2018; Algabbani and Algabbani,, 2020). While researchers have extensively documented this problem through surveys and interviews (Boratas and Kilic,, 2018; Fernandez-Lazaro et al.,, 2019; Algabbani and Algabbani,, 2020; Najjuma et al.,, 2020; Schober et al.,, 2021), the studies—and ultimately understanding of treatment non-adherence—remain limited by small sample sizes and self-reporting bias (Adams et al.,, 1999; Stirratt et al.,, 2015). Physical solutions to monitor and encourage adherence such as electronic pill caps have shown promise in controlled settings but remain impractical for large-scale deployment due to high costs and implementation challenges (Parker et al.,, 2007; Mauro et al.,, 2019).
@@
These measurement challenges take on new urgency as healthcare systems increasingly rely on machine learning (ML) models trained on electronic health records (EHRs) to guide treatment decisions (Komorowski et al.,, 2018; Brugnara et al.,, 2020; Zheng et al.,, 2021; Mroz et al.,, 2024; Yi et al.,, 2024; Shen et al.,, 2024; Chen et al.,, 2022). These machine learning models learn from historical patient data, which assume that prescribed treatments were actually taken. However, this introduces an implicit bias—models trained on non-adherent patients learn patterns that misrepresent true treatment effects. This implicit bias may degrade model performance and disproportionately impact underserved populations, who often face greater barriers to treatment adherence (Bosworth et al.,, 2006; Schober et al.,, 2021).
@@
Recent advances in large language models (LLMs) have shown that LLMs can advance medical understanding by accurately extracting information from EHRs (Agrawal et al.,, 2022; Goel et al.,, 2023). Instead of relying on self-reported treatment adherence from questionnaires and interviews, LLMs could serve as a powerful tool for identifying treatment non-adherence directly from EHRs. By analyzing rich but unstructured clinical notes, LLMs can detect documented instances of missed medications, unfilled prescriptions, and patient-reported barriers to adherence, enabling systematic assessment of treatment non-adherence across large patient populations.
@@
In this study, we examine hypertension treatment non-adherence using EHR data from a large academic hospital by leveraging an LLM to analyze clinical notes, and further investigate its impact on causal inference and ML model performance (Figure 1). With a cohort of 3,623 patients, we identify 786 (21.7%) cases of non-adherence and extract demographic and clinical factors that are statistically significant. Additionally, we apply topic modeling to clinical notes revealing underlying reasons for non-adherence.
@@ 2.1 Treatment adherence analysis in hypertension
Multiple studies have investigated treatment adherence among patients with hypertension (Boratas and Kilic,, 2018; Uchmanowicz et al.,, 2018; Algabbani and Algabbani,, 2020; Najjuma et al.,, 2020; Schober et al.,, 2021). These studies are mainly cross-sectional, with a cohort of admitted patients collected at a fixed time point, and treatment adherence is typically measured through questionnaires and interviews. For instance, Algabbani and Algabbani, (2020) conducted a study in Saudi Arabia involving 306 hypertensive outpatients, finding that only 42.2% of participants adhered to their antihypertensive medications. Boratas and Kilic, (2018) conducted a similar study of 147 hypertensive patients, identifying factors such as age and duration of hypertension to be significant. However, due to their reliance on questionnaire and interview data, they often have small sample sizes (e.g., less than 300 patients) and self-reporting bias (Adams et al.,, 1999; Stirratt et al.,, 2015), which limits their representativeness and can even lead to contradictory conclusions. In contrast, our work conducts the first large-scale analysis utilizing EHR, with a significantly larger sample size of 3,623 patients.
@@ 2.2 Machine learning and treatment adherence
Machine learning has been used to identify individual risk factors associated with treatment non-adherence (Koesmahargyo et al.,, 2020; Gichuhi et al.,, 2023; Burgess-Hull et al.,, 2023). Gichuhi et al., (2023) developed ML algorithms and found SVM achieved 91.28% accuracy in predicting tuberculosis treatment non-adherence. Instead of predicting treatment adherence, our work focuses on analyzing the impact of treatment non-adherence bias on downstream model performance. Other studies have applied natural language processing (NLP) to analyze surveys to better understand treatment non-adherence (Anglin et al.,, 2021; Lin et al.,, 2022; Chan et al.,, 2024). Chan et al., (2024) applied NLP to free-text responses from questionnaires completed by type 2 diabetes patients, identifying key reasons for non-adherence. Unlike questionnaires, our work leverages treatment adherence information extracted from clinical notes using LLMs. Lastly, Zhong et al., (2022) applied ML while accounting for adherence information when analyzing treatment effects in a randomized controlled trial. To our knowledge, our study is the first to leverage LLMs for extracting treatment adherence information from clinical notes and evaluating its impact on downstream causal inference and predictive model performance.
@@ 3.1 Hypertension cohort selection
Our analysis focuses on ten commonly prescribed hypertension medications: amlodipine, losartan, lisinopril, benazepril, carvedilol, hydralazine, hydrochlorothiazide, clonidine, spironolactone, and metoprolol (hea, 2024a, ). Therefore, we excluded pairs in which the first visit lacked a medication record on this list, as well as pairs with missing or invalid notes during the second visit. We further focus on pairs where the interval between visits is between one month and one year. Lastly, we filtered out patients with unknown demographic information for the purpose of analysis. This resulted in a final cohort of 3,623 patients with 5,952 visit pairs. The cohort selection process is summarized in Appendix A.
Demographic information, including sex, age, race, and marital status, was extracted from patient records. Four clinical factors were further derived from the EHR, many of which have been shown to be associated with hypertension non-adherence (Boratas and Kilic,, 2018; Algabbani and Algabbani,, 2020). These factors include the duration between the two visits in the pair, the duration of hypertension, the number of primary care visits and the number of comorbidities. We quantified comorbidities using the Charlson Comorbidity Index (CCI) (Charlson et al.,, 1987) and the Elixhauser Comorbidity Index (ECI) (Elixhauser et al.,, 1998), which condensed diagnoses into 17 and 31 well-defined comorbidity categories respectively. The demographic and clinical characteristics of the selected cohort are summarized in Table 1. We detail the comorbidity categories along with other features used in the study in Appendix B.
@@ 3.2 LLM configuration and prompt engineering
We used the GPT-4o model (OpenAI et al.,, 2024) (version 2024-05-13) via the HIPAA-compliant Microsoft Azure API, with the temperature set to 0 and all other parameters left at default. For each pair of visits, we provided the prescription record from the first visit and the clinical notes from the second visit to the model to assess adherence to the prescribed medication.
@@ 4.1 Factors associated with treatment non-adherence
To identify factors associated with nonadherence to treatment, we begin by performing unadjusted logistic regression, with the results presented in Table 1. We find three factors that are statistically significant ($p<0.05$): age ($p=0.036$), one-hot encoding for Black race ($p=0.023$), and the number of comorbidities ($p=0.045$).
Our findings indicate that younger patients are less likely to adhere to treatment, aligning with previous research (Boratas and Kilic,, 2018) that suggests adherence improves with age as patients become more accustomed to managing their diagnoses. Additionally, we find that Black patients exhibit higher rates of non-adherence. In the U.S., Black individuals have a higher prevalence of uncontrolled hypertension than White individuals (Aggarwal et al.,, 2021), and our finding further highlighting the need for greater attention to prevent further exacerbation of racial disparities in hypertension control. Lastly, a lower number of comorbidities is associated with a higher rate of non-adherence, possibly because patients with fewer health conditions may perceive their treatment as less essential.
@@ 4.2 Uncovering reasons for treatment non-adherence
We employ Bidirectional Encoder Representations from Transformers (BERT)-based topic modeling (Grootendorst,, 2022) on the extracted non-adherent clinical notes to uncover reasons. This utilizes a BERT architecture (Devlin et al.,, 2019) to generate embeddings from the extracted notes and applies Uniform Manifold Approximation and Projection (UMAP) (McInnes et al.,, 2020) to reduce the dimensionality of the embeddings. The reduced embeddings are subsequently clustered using HDBSCAN (Campello et al.,, 2013), and key terms for each cluster are identified using class-based Term Frequency-Inverse Document Frequency (c-TF-IDF).
@@ 5.2 Causal inference for treatment effect estimation
Amlodipine and Lisinopril are the most commonly prescribed medications for hypertension, leading to numerous randomized controlled trials (RCTs) comparing their treatment effects (Cappuccio et al.,, 1993; Naidu et al.,, 2000). However, due to the high cost of RCTs, various causal inference methods have been developed to estimate treatment effects from observational data (Pearl,, 2009; Austin,, 2011; Shalit et al.,, 2017; Künzel et al.,, 2019). Among them, Inverse Probability Weighting (IPW) is one of the most widely used methods (Austin,, 2011), providing an unbiased estimation of the Average Treatment Effect (ATE) by adjusting for confounding. We start by demonstrating the impact of treatment non-adherence bias on ATE estimation using IPW.
@@
Results. The results are presented in Table 4. Without filtering for non-adherent data, amlodipine lowers diastolic blood pressure by 1.75 mmHg but increases systolic blood pressure by 0.06 mm Hg compared to lisinopril. After excluding non-adherent data, amlodipine lowers diastolic blood pressure by 1.40 mmHg and also reduces systolic blood pressure by 0.11 mmHg compared to lisinopril. This result shows a reversal in the estimated treatment effect for systolic blood pressure reduction before and after excluding non-adherent data.
Table 4: Estimated ATE of medication on blood pressure reduction using IPW. Notably, excluding non-adherent data reverses the conclusion on the treatment effect of systolic blood pressure reduction, causing the estimate to flip from -0.06 to 0.11 mmHg.
Dataset Blood Pressure Reduction
Diastolic Systolic
Full Dataset 1.75 -0.06
Adherent Data Only 1.40 0.11
@@ 5.3 Supervised learning for treatment outcome prediction
We now demonstrate the impact of treatment non-adherence bias on predictive modeling performance. Following a common setup in the literature (Mroz et al.,, 2024; Yi et al.,, 2024), we use patients’ EHR data with treatment prescriptions and blood pressure measurements from their first visit as covariates. The target to predict is whether the blood pressure will be normal at their second visit. Following the guidelines of the American Heart Association (hea, 2024b, ), we define normal blood pressure as having a systolic value of less than 120 and a diastolic value of less than 80. To evaluate model performance in predicting outcomes, we use 500 adherent samples as the test set in all subsequent experiments. We test exclusively on adherent patients since our goal is to evaluate model performance in scenarios where treatments are followed as prescribed, representing the intended clinical use case. A detailed description of the features used is provided in Appendix B.
@@ 5.3.1 Effect of varying treatment non-adherence data ratios on model performance and fairness
Experiment Setup. We fix the training set size at 300 and vary the proportion of non-adherent data in the training set across $\{0\%,10\%,30\%,50\%,70\%,90\%\}$ to evaluate the impact of treatment non-adherence bias on model performance. We train logistic regression and random forest models, both commonly used for modeling tabular EHR data and assess performance using AUROC on the test set. Beyond performance, ensuring fair decision-making is also a critical consideration in healthcare. Let $A$ denote the sensitive attribute (e.g., race), $Y$ represent the true outcome, and $\hat{Y}$ denote the predicted outcome. Demographic parity (Dwork et al.,, 2012) difference measures the disparity in the likelihood of receiving a positive prediction between groups, i.e.,
@@
Equal odds (Hardt et al.,, 2016) difference compares both true positive rates and false positive rates across groups, i.e.,

### v2

# Revealing Treatment Non‑Adherence Bias in Clinical Machine Learning Using Large Language Models

@@ Abstract
Machine learning systems trained on electronic health records (EHRs) increasingly guide treatment decisions, but their reliability depends on the critical assumption that patients follow the prescribed treatments recorded in EHRs. Using EHR data from 3,623 hypertension patients, we investigate how treatment non-adherence introduces implicit bias that can fundamentally distort both causal inference and predictive modeling. By extracting patient adherence information from clinical notes using a large language model (LLM), we identify 786 patients (21.7%) with medication non-adherence. We further uncover key demographic and clinical factors associated with non-adherence, as well as patient-reported reasons including side effects and difficulties obtaining refills. Our findings demonstrate that this implicit bias can not only reverse estimated treatment effects, but also degrade model performance by up to 5% while disproportionately affecting vulnerable populations by exacerbating disparities in decision outcomes and model error rates. This highlights the importance of accounting for treatment non-adherence in developing responsible and equitable clinical machine learning systems.
Data and Code Availability
Our study utilizes EHR data from UCSF. Details on the data and preprocessing steps are provided in the following sections. While we will make the code publicly available, the data cannot be shared due to data use agreement.
Institutional Review Board (IRB)
Our work was reviewed by an IRB and declared exempt as it focuses on retrospective analysis using de-identified data.
@@ 1 Introduction
Treatment non-adherence is a pervasive and persistent challenge in healthcare. Researchers estimate that poor medication adherence leads to 125,000 preventable deaths annually in the U.S. and contributes to \$100-\$300 billion in avoidable healthcare costs (Benjamin, 2012). This issue is particularly prevalent among patients with chronic conditions such as hypertension, with 40-50% failing to take their medications as prescribed (Kleinsinger, 2018; Algabbani and Algabbani, 2020). While researchers have extensively documented this problem through surveys and interviews (Boratas and Kilic, 2018; Fernandez-Lazaro et al., 2019; Algabbani and Algabbani, 2020; Najjuma et al., 2020; Schober et al., 2021), the studies—and ultimately understanding of treatment non-adherence—remain limited by small sample sizes and self-reporting bias (Adams et al., 1999; Stirratt et al., 2015). Physical solutions to monitor and encourage adherence such as electronic pill caps have shown promise in controlled settings but remain impractical for large-scale deployment due to high costs and implementation challenges (Parker et al., 2007; Mauro et al., 2019).
@@
These measurement challenges take on new urgency as healthcare systems increasingly rely on machine learning (ML) models trained on electronic health records (EHRs) to guide treatment decisions (Komorowski et al., 2018; Brugnara et al., 2020; Zheng et al., 2021; Mroz et al., 2024; Yi et al., 2024; Shen et al., 2024; Chen et al., 2022). These machine learning models learn from historical patient data, which assume that prescribed treatments were actually taken. However, this introduces an implicit bias—models trained on non-adherent patients learn patterns that misrepresent true treatment effects. This implicit bias may degrade model performance and disproportionately impact underserved populations, who often face greater barriers to treatment adherence (Bosworth et al., 2006; Schober et al., 2021).
@@
Recent advances in large language models (LLMs) have shown that LLMs can advance medical understanding by accurately extracting information from EHRs (Agrawal et al., 2022; Goel et al., 2023). Instead of relying on self-reported treatment adherence from questionnaires and interviews, LLMs could serve as a powerful tool for identifying treatment non-adherence directly from EHRs. By analyzing rich but unstructured clinical notes, LLMs can detect documented instances of missed medications, unfilled prescriptions, and patient-reported barriers to adherence, enabling systematic assessment of treatment non-adherence across large patient populations.
@@
In this study, we examine hypertension treatment non-adherence using EHR data from UCSF by leveraging an LLM to analyze clinical notes, and further investigate its impact on causal inference and ML model performance (Figure 1). With a cohort of 3,623 patients, we identify 786 (21.7%) cases of non-adherence and extract demographic and clinical factors that are statistically significant. Additionally, we apply topic modeling to clinical notes revealing underlying reasons for non-adherence.
@@ 2.1 Treatment adherence analysis in hypertension
Multiple studies have investigated treatment adherence among patients with hypertension (Boratas and Kilic, 2018; Uchmanowicz et al., 2018; Algabbani and Algabbani, 2020; Najjuma et al., 2020; Schober et al., 2021). These studies are mainly cross-sectional, with a cohort of admitted patients collected at a fixed time point, and treatment adherence is typically measured through questionnaires and interviews. For instance, Algabbani and Algabbani (2020) conducted a study in Saudi Arabia involving 306 hypertensive outpatients, finding that only 42.2% of participants adhered to their antihypertensive medications. Boratas and Kilic (2018) conducted a similar study of 147 hypertensive patients, identifying factors such as age and duration of hypertension to be significant. However, due to their reliance on questionnaire and interview data, they often have small sample sizes (e.g., less than 300 patients) and self-reporting bias (Adams et al., 1999; Stirratt et al., 2015), which limits their representativeness and can even lead to contradictory conclusions. In contrast, our work conducts the first large-scale analysis utilizing EHR, with a significantly larger sample size of 3,623 patients.
@@ 2.2 Machine learning and treatment adherence
Machine learning has been used to identify individual risk factors associated with treatment non-adherence (Koesmahargyo et al., 2020; Gichuhi et al., 2023; Burgess-Hull et al., 2023). Gichuhi et al. (2023) developed ML algorithms and found SVM achieved 91.28% accuracy in predicting tuberculosis treatment non-adherence. Instead of predicting treatment adherence, our work focuses on analyzing the impact of treatment non-adherence bias on downstream model performance. Other studies have applied natural language processing (NLP) to analyze surveys to better understand treatment non-adherence (Anglin et al., 2021; Lin et al., 2022; Chan et al., 2024). Chan et al. (2024) applied NLP to free-text responses from questionnaires completed by type 2 diabetes patients, identifying key reasons for non-adherence. Unlike questionnaires, our work leverages treatment adherence information extracted from clinical notes using LLMs. Lastly, Zhong et al. (2022) applied ML while accounting for adherence information when analyzing treatment effects in a randomized controlled trial. To our knowledge, our study is the first to leverage LLMs for extracting treatment adherence information from clinical notes and evaluating its impact on downstream causal inference and predictive model performance.
@@ 3.1 Hypertension cohort selection
Our analysis focuses on ten commonly prescribed hypertension medications: amlodipine, losartan, lisinopril, benazepril, carvedilol, hydralazine, hydrochlorothiazide, clonidine, spironolactone, and metoprolol (hea, 2024a). Therefore, we excluded pairs in which the first visit lacked a medication record on this list, as well as pairs with missing or invalid notes during the second visit. We further focus on pairs where the interval between visits is between one month and one year. Lastly, we filtered out patients with unknown demographic information for the purpose of analysis. This resulted in a final cohort of 3,623 patients with 5,952 visit pairs. The cohort selection process is summarized in Appendix A.
Demographic information, including sex, age, race, and marital status, was extracted from patient records. Four clinical factors were further derived from the EHR, many of which have been shown to be associated with hypertension non-adherence (Boratas and Kilic, 2018; Algabbani and Algabbani, 2020). These factors include the duration between the two visits in the pair, the duration of hypertension, the number of primary care visits and the number of comorbidities. We quantified comorbidities using the Charlson Comorbidity Index (CCI) (Charlson et al., 1987) and the Elixhauser Comorbidity Index (ECI) (Elixhauser et al., 1998), which condensed diagnoses into 17 and 31 well-defined comorbidity categories respectively. The demographic and clinical characteristics of the selected cohort are summarized in Table 1. We detail the comorbidity categories along with other features used in the study in Appendix B.
@@ 3.2 LLM configuration and prompt engineering
We used the GPT-4o model (OpenAI et al., 2024) (version 2024-05-13) via the HIPAA-compliant Microsoft Azure API, with the temperature set to 0 and all other parameters left at default. For each pair of visits, we provided the prescription record from the first visit and the clinical notes from the second visit to the model to assess adherence to the prescribed medication.
@@ 4.1 Factors associated with treatment non-adherence
To identify factors associated with nonadherence to treatment, we begin by performing unadjusted logistic regression, with the results including confidence intervals and p-values presented in Table 1. We find three factors that are statistically significant ($p<0.05$): age ($p=0.036$), one-hot encoding for Black race ($p=0.023$), and the number of comorbidities ($p=0.045$).
Our findings indicate that younger patients are less likely to adhere to treatment, aligning with previous research (Boratas and Kilic, 2018) that suggests adherence improves with age as patients become more accustomed to managing their diagnoses. Additionally, we find that Black patients exhibit higher rates of non-adherence. In the U.S., Black individuals have a higher prevalence of uncontrolled hypertension than White individuals (Aggarwal et al., 2021), and our finding further highlighting the need for greater attention to prevent further exacerbation of racial disparities in hypertension control. Lastly, a lower number of comorbidities is associated with a higher rate of non-adherence, possibly because patients with fewer health conditions may perceive their treatment as less essential.
@@ 4.2 Uncovering reasons for treatment non-adherence
We employ Bidirectional Encoder Representations from Transformers (BERT)-based topic modeling (Grootendorst, 2022) on the extracted non-adherent clinical notes to uncover reasons. This utilizes a BERT architecture (Devlin et al., 2019) to generate embeddings from the extracted notes and applies Uniform Manifold Approximation and Projection (UMAP) (McInnes et al., 2020) to reduce the dimensionality of the embeddings. The reduced embeddings are subsequently clustered using HDBSCAN (Campello et al., 2013), and key terms for each cluster are identified using class-based Term Frequency-Inverse Document Frequency (c-TF-IDF).
@@ 5.1 Treatment non-adherence leads to worse blood pressure outcomes
Table 4: Estimated ATE of medication on blood pressure reduction using different causal inference methods. Notably, excluding non-adherent data leads to a lower estimated effect on the diastolic blood pressure reduction and reverses the conclusion on the treatment effect of systolic blood pressure reduction consistently across methods.
Diastolic Reduction (mmHg) Systolic Reduction (mmHg)
IPW S-Learner T-Learner X-Learner IPW S-Learner T-Learner X-Learner
Full Dataset 1.75 0.77 1.44 1.51 -0.06 -0.05 -0.12 -0.14
Adherent Data Only 1.40 0.57 0.97 0.92 0.11 0.08 0.06 0.07
@@ 5.2 Causal inference for treatment effect estimation
Amlodipine and Lisinopril are the most commonly prescribed medications for hypertension, leading to numerous randomized controlled trials (RCTs) comparing their treatment effects (Cappuccio et al., 1993; Naidu et al., 2000). However, due to the high cost of RCTs, various causal inference methods have been developed to estimate treatment effects from observational data (Pearl, 2009; Austin, 2011; Shalit et al., 2017; Künzel et al., 2019). Among them, Inverse Probability Weighting (IPW) provides an unbiased estimation of the Average Treatment Effect (ATE) by adjusting for confounding (Austin, 2011). Meta-learners such as the S-learner, T-learner and X-learner have been proposed to improve treatment effect estimation by leveraging flexible ML models to learn heterogeneous treatment effects (Künzel et al., 2019). We start by demonstrating the impact of treatment non-adherence bias on ATE estimation using IPW and meta-learners.
@@
Results. The results are presented in Table 4. Across all models, excluding non-adherent data leads to a lower estimated effect on the diastolic blood pressure reduction and a reversal in the systolic blood pressure reduction. Specifically when using IPW, without filtering for non-adherent data, amlodipine lowers diastolic blood pressure by 1.75 mmHg but increases systolic blood pressure by 0.06 mm Hg compared to lisinopril. After excluding non-adherent data, amlodipine lowers diastolic blood pressure by 1.40 mmHg and also reduces systolic blood pressure by 0.11 mmHg compared to lisinopril. This result shows a reversal in the estimated treatment effect for systolic blood pressure reduction before and after excluding non-adherent data. The same reversal trend also holds across different meta-learners (e.g., using the X-learner on the full dataset, amlodipine reduces systolic blood pressure by 0.14 mmHg compared to lisinopril, whereas after excluding non-adherent data, the X-learner estimates an increase of 0.07 mmHg in systolic blood pressure reduction). The results demonstrate that non-adherence bias can lead to systematic distortions in treatment effect estimation.
@@ 5.3 Supervised learning for treatment outcome prediction
We now demonstrate the impact of treatment non-adherence bias on predictive modeling performance. Following a common setup in the literature (Mroz et al., 2024; Yi et al., 2024), we use patients’ EHR data with treatment prescriptions and blood pressure measurements from their first visit as covariates. The target to predict is whether the blood pressure will be normal at their second visit. Following the guidelines of the American Heart Association (hea, 2024b), we define normal blood pressure as having a systolic value of less than 120 and a diastolic value of less than 80. To evaluate model performance in predicting outcomes, we use 500 adherent samples as the test set in all subsequent experiments. We test exclusively on adherent patients since our goal is to evaluate model performance in scenarios where treatments are followed as prescribed, representing the intended clinical use case. A detailed description of the features used is provided in Appendix B.
@@ 5.3.1 Effect of varying treatment non-adherence data ratios on model performance and fairness
Experiment Setup. We fix the training set size at 300 and vary the proportion of non-adherent data in the training set across $\{0\%,10\%,30\%,50\%,70\%,90\%\}$ to evaluate the impact of treatment non-adherence bias on model performance. We train logistic regression and random forest models, both commonly used for modeling tabular EHR data and assess performance using AUROC on the test set. Beyond performance, ensuring fair decision-making is also a critical consideration in healthcare. Let $A$ denote the sensitive attribute (e.g., race), $Y$ represent the true outcome, and $\hat{Y}$ denote the predicted outcome. Demographic parity (Dwork et al., 2012) difference measures the disparity in the likelihood of receiving a positive prediction between groups, i.e.,
@@
Equal odds (Hardt et al., 2016) difference compares both true positive rates and false positive rates across groups, i.e.,